# Mean-field underdamped Langevin dynamics and its space-time discretization

### 01 Jan 2024 v2

@@ Abstract
We propose a new method called the N-particle underdamped Langevin algorithm for optimizing a special class of non-linear functionals defined over the space of probability measures. Examples of problems with this formulation include training neural networks in the mean-field regime, density estimation, and kernel Stein discrepancy minimization. Our algorithm is based on a novel space-time discretization of the mean-field underdamped Langevin dynamics, for which we provide a new, fast mixing guarantee. In addition, we demonstrate that our algorithm converges globally in total variation distance, bridging the theoretical gap between the dynamics and its practical implementation.
Qiang ${}^{\star}$ fuqiang7@mail2.sysu.edu.cn Ashia ${}^{\dagger}$ ashia07@mit.edu
@@ 1 Introduction
Many problems in data science can be posed as an entropy-regularized mean-field optimization (EMO) problem, described by
where $\lambda>0$; $\mathcal{P}_{2}(\mathbb{R}^{d})$ is the measure space on $\mathbb{R}^{d}$ with finite second moment; $F$ is a non-linear convex functional and $\text{Ent}(\mu)=\int\mu\log\mu$ is the negative entropy of $\mu$. Classical applications include training neural networks in the mean-field regime (Mei et al., 2018; Hu et al., 2019; Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022; Claisse et al., 2023; Suzuki et al., 2023), density estimation via maximum mean discrepancy minimization (Gretton et al., 2006; Arbel et al., 2019; Chizat, 2022; Suzuki et al., 2023) and kernel Stein discrepancy minimization (Liu et al., 2016; Chwialkowski et al., 2016; Suzuki et al., 2023).
@@
A popular approach for finding a minimizer of an EMO problem is based on the mean-field Langevin dynamics. When the problem is convex, Hu et al. (2019) show the mean-field Langevin dynamics convergences to the minimizer asymptotically; and when the problem satisfies a uniform logarithmic Sobolev inequality, several works have established an exponentially fast convergence (Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022). The mean-field Langevin dynamics, however, cannot be implemented and bridging this gap requires both space and time discretizations of the dynamics. A time-discretization of the mean-field Langevin dynamics was analyzed by Nitanda et al. (2022) who extend the interpolation argument introduced by Vempala and Wibisono (2019) to a non-linear Fokker-Planck equation. They establish the non-asymptotic convergence of the time-discretized dynamics in the energy gap. A space-discretization was studied by Chen et al. (2022) who show that the finite-particle approximation to the density of the mean-field Langevin dynamics (referred to as a uniform-in-time propagation of chaos) converges exponentially fast, with a bias related to the number of particles. More practically, Suzuki et al. (2023) analyze a space-time discretization of the mean-field Langevin dynamics and establish the non-asymptotic convergence of the resulting algorithm to a biased limit related to both the particles and stepsize. Their analysis applies to several important learning problems and improves the results of the standard gradient Langevin dynamics.
@@
A natural candidate method for finding solutions to EMO problems faster is the mean-field underdamped Langevin dynamics. These dynamics resemble several techniques for adding momentum to gradient descent in optimization, many of which are known to result in provable faster convergence in a variety of settings (Nesterov, 1983; Wilson et al., 2016; Laborde and Oberman, 2020; Fu et al., 2023). Moreover, training neural networks using momentum-based gradient descent is considered effective in several applications (Sutskever et al., 2013; Kingma and Ba, 2014; Ruder, 2016). Kazeykina et al. (2020) and Chen et al. (2023) confirm that even naive space-time discretizations of the mean-field underdamped Langevin dynamics have impressive empirical performance when compared to various proposed discretizations of the mean-field Langevin dynamics on applications such as training of neural networks.
@@
In an effort to better understand the behavior of the mean-field underdamped Langevin dynamics, Chen et al. (2023) establish an exponential rate convergence of it and its finite-particle system under assumptions that are easy to verify for specific objectives of training neural networks. In addition, Chen et al. (2023) implement an Euler-Maruyama discretization of the finite-particle system and show that it performs empirically faster when compared with the space-time discretization of the mean-field Langevin dynamics in a variety of settings, such as training a toy neural network model. However, space-time discretizations of the mean-field underdamped Langevin dynamics are not yet theoretically well understood. Furthermore, the rate obtained by Chen et al. (2023) for the dynamics does not resemble an accelerated rate when compared with Nitanda et al. (2022) and Chen et al. (2022) who analyze the overdamped setting.
@@
Despite recent results, a remaining question is whether we can theoretically characterize the behavior of an implementable algorithm based on discretizing the mean-field underdamped dynamics. If there is a limiting bias, how does it scale with the number of particles and other problem parameters? Ideally, this characterization would give a sharper rate of convergence than Suzuki et al. (2023)’s space-time discretization of the mean-field Langevin dynamics, suggesting there might be a worst-case advantage to adding momentum in the mean-field setting. In this paper, we introduce a fast implementable algorithm for solving EMO problems based on the mean-field underdamped Langevin dynamics. We prove that our proposed algorithm converges to a small limiting bias under a set of benign assumptions. In particular, our contributions are summarized as follows.
@@
    We sharpen the convergence guarantees for both the mean-field underdamped Langevin dynamics and its space-discretization under the same set of assumptions utilized by Chen et al. (2023) (summarized in Theorems 3.1 and 3.2). In particular, we sharpen the dependence on the smoothness constant.
@@
    We introduce a novel space-time discretization of the mean-field underdamped Langevin dynamics which we call the N-particle underdamped Langevin algorithm (Algorithm 1). We show the global convergence of our discretization in terms of the total variation (TV) distance under additional assumptions. Importantly, our results improve on Suzuki et al. (2023)’s analysis of the space-time discretization of the mean-field Langevin dynamics, and our additional assumptions are satisfied by the same examples introduced by Suzuki et al. (2023); these results provide theoretical characterizations of momentum methods in several real-world applications including training neural networks and density estimation.
@@
The remainder of this work is organized as follows. Section 2 presents the formal definitions and assumptions as well as related works. Section 3 proposes our main methods and theoretical results. Section 4 discusses the application of our methods to some classical problems. Section 5 describes our numerical experiments verifying the effectiveness of our proposed methods.
@@ 2.2 Background
The gradient flow structure in 2-Wasserstein metric of the EMO functional is called the mean-field Langevin dynamics,
@@
Hu et al. (2019) show that the minimizer of MLD, $\mu_{*}(x)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu_{*},x)\right),$ corresponds to solutions of an EMO problem under mild conditions. We study the mean-field underdamped Langevin dynamics,
@@
and show a new sharp mixing-time bound; here, $\gamma>0$ is the damping coefficient; $\mu_{t}(x,v)=\text{Law}(x_{t},v_{t})$ and $\mu_{t}^{x}\triangleq\text{Law}(x_{t})=\int\mu_{t}(x,v)\mathrm{d}v$ is the $x$-marginal of $\mu_{t}$. The limiting distribution of MULD is the solution to,
@@
where a momentum term is added to the EMO problem. To obtain the solution of the EMO problem, it suffices to marginalize the minimizer of (1) given by, $\mu_{*}(x,v)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2\lambda}\|v\|^{2}\right).$ This work also sharpens the analysis of the space-discretization of MULD introduced by Chen et al. (2023), which we refer to as the $N$-particle underdamped Langevin dynamics,
@@
$\mathrm{d}v_{t}^{i}$ $=-\gamma v_{t}^{i}\mathrm{d}t\vspace{-10cm}-D_{\mu}F(\mu_{\textbf{x}_{t}},x^{i}_{t})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}\mathrm{B}^{i}_{t};$
$i=1,...,N$ where $\mu_{\textbf{x}}\triangleq\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$; $\mu^{i}\triangleq\text{Law}(x^{i},v^{i})$ and $(\mathrm{B}_{t}^{i})_{i=1}^{N}$ are $d$-dimensional Brownian motions independent of $i$. While interesting, a time-discretization of NULD is necessary to run the method on a machine. We explore two time-discretization techniques in this paper.
To motivate our algorithm as a time-discretization of NULD, we review discretizations of the underdamped Langevin dynamics (ULD), which is a special case of MULD where $F(\mu)=\int V(x)\mu(dx)$ is a linear functional of $\mu$:
@@
for stepsize $h$, $\xi_{k}\sim\mathcal{N}(0,I_{d})$ and $t\in[kh,(k+1)h]$, has been well-studied and shown to incur the largest error (Kloeden et al., 1995; Platen and Bruti-Liberati, 2010). Recently, however, several works derive the convergence of the ULMC in Wasserstein distance (Cheng et al., 2018), KL divergence (Ma et al., 2021) and Rényi divergence (Zhang et al., 2023) using a more precise discretization scheme, often referred to as the left point method (LPM) (Hu et al., 2021) and given by
@@
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-\nabla V(x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}\mathrm{B}_{t},$
@@ 2.3 Definitions and assumptions
Note that if $\hat{\mu}^{x}(x)\propto\exp(-\frac{\delta F}{\delta\mu}(\mu^{x},x))$ satisfies LSI with constant $\rho>0$, Assumption 2.3 is satisfied with the choice $\mathscr{C}_{\textsf{LSI}}=\frac{1}{2}\land\rho$. And we refer our readers to Chen et al. (2023) and Suzuki et al. (2023) for the verification of Assumptions 2.1-2.3 in a variety of settings, including the neural network setting and density estimation setting. Beyond Assumptions 2.1-2.3, we introduce four additional assumptions that will be required to show the global convergence of our space-time discretization.
@@
Notably, Suzuki et al. (2023) assume that $F$ can be decomposed as $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ where $\|D_{\mu}U(\mu,x)\|\leq R$ for any $\mu\in\mathcal{P}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, and where $r(x)$ is $\lambda_{2}$-smooth with $\nabla r(0)=0$ in order to establish the convergence of their space-time discretization of MLD. Thus, their assumption, $\|D_{\mu}F(\mu,x)\|\leq\|D_{\mu}U(\mu,x)\|+\|\nabla r(x)\|\leq R+\lambda_{2}\|x\|$, implies Assumption 2.4 with the choice $\mathscr{L}\geq\max\{R,\lambda_{2}\}$. The next three assumptions are needed for bounding the second moment of the iterates $(x_{t},v_{t})$ and $(x_{t}^{i},v_{t}^{i})$ along MULD and NULD, which is crucial for the establishment of our discrete-time convergence:
@@ 3 N-particle underdamped Langevin algorithm
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-D_{\mu}F(\mu^{x}_{k},x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}\mathrm{B}_{t},$
@@
$\mathrm{d}v_{t}^{i}$ $=\gamma v^{i}_{t}\mathrm{d}t-D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}\mathrm{B}^{i}_{t}$
@@ 3.1 Convergence analysis
Space-time discretizations NLA (Suzuki et al., 2023) ${\Theta}\left(\kappa\mathscr{L}^{3}d/\epsilon^{2}\right)$ $\widetilde{\Theta}\left(\kappa^{2}\mathscr{L}d/\epsilon^{2}\right)$
@@ 4 Applications of Algorithm 1
In Section 1, we briefly mention three examples on the applications of problem (1). In this section, we will show how Algorithm 1 can be applied to training mean-field two-layer neural networks and density estimation via maximum mean discrepancy (MMD) minimization.

### 17 Jan 2024 v3

# Mean-field underdamped Langevin dynamics and its spacetime discretization

@@ Abstract
We propose a new method called the N-particle underdamped Langevin algorithm for optimizing a special class of non-linear functionals defined over the space of probability measures. Examples of problems with this formulation include training mean-field neural networks, maximum mean discrepancy minimization and kernel Stein discrepancy minimization. Our algorithm is based on a novel spacetime discretization of the mean-field underdamped Langevin dynamics, for which we provide a new, fast mixing guarantee. In addition, we demonstrate that our algorithm converges globally in total variation distance, bridging the theoretical gap between the dynamics and its practical implementation.
Qiang ${}^{\star}$ fuqiang7@mail2.sysu.edu.cn Ashia ${}^{\dagger}$ ashia@mit.edu
@@ 1 Introduction
Many problems in data science can be posed as an entropy-regularized mean-field optimization (EMO) problem. This includes classical applications such as training mean-field neural networks (Mei et al., 2018; Hu et al., 2019; Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022; Claisse et al., 2023; Suzuki et al., 2023), density estimation via maximum mean discrepancy (MMD) minimization (Gretton et al., 2006; Arbel et al., 2019; Chizat, 2022; Suzuki et al., 2023) and sampling via kernel Stein discrepancy (KSD) minimization (Liu et al., 2016; Chwialkowski et al., 2016; Suzuki et al., 2023). A popular approach for solving an EMO problem is the mean-field Langevin dynamics. When the problem can be expressed as optimizing a convex functional, Hu et al. (2019) have demonstrated that mean-field Langevin dynamics converge to a minimizer asymptotically. Additionally, if the problem satisfies a uniform logarithmic Sobolev inequality, several studies have established that this convergence occurs exponentially quickly (Chizat, 2022; Nitanda et al., 2022; Chen et al., 2022). However, implementing the mean-field Langevin dynamics is not a straightforward task; to arrive at a practical algorithm, both spatial and temporal discretizations of the dynamics is required.
@@
Towards this end, a time-discretization of the mean-field Langevin dynamics was analyzed by Nitanda et al. (2022) who extend the interpolation argument introduced by Vempala and Wibisono (2019) to a non-linear Fokker-Planck equation. They establish the non-asymptotic convergence of the discrete-time process in the energy gap. A space-discretization was studied by Chen et al. (2022) who show that the finite-particle approximation to the density of the mean-field Langevin dynamics (referred to as a uniform-in-time propagation of chaos) converges to the minimizer of EMO problem exponentially fast, with a bias related to the number of particles. More practically, Suzuki et al. (2023) analyze a spacetime discretization of the mean-field Langevin dynamics and establish the non-asymptotic convergence of the resulting algorithm to a biased limit related to both the number of particles used and stepsize. Their analysis applies to several important learning problems and improves the results of the standard gradient Langevin dynamics.
@@
A natural candidate method for finding solutions to EMO problems faster is the mean-field underdamped Langevin dynamics. These dynamics resemble several techniques for adding momentum to gradient descent in optimization, many of which are known to result in provably faster convergence in a variety of settings (Nesterov, 1983; Wilson et al., 2016; Laborde and Oberman, 2020; Fu et al., 2023). Moreover, training neural networks using momentum-based gradient descent is considered effective in several applications (Sutskever et al., 2013; Kingma and Ba, 2014; Ruder, 2016). Kazeykina et al. (2020) and Chen et al. (2023) confirm that the naive spacetime discretization of the mean-field underdamped Langevin dynamics have impressive empirical performance when compared to naive discretization of the mean-field Langevin dynamics on applications such as training mean-field neural networks.
@@
In an effort to better understand the behavior of the mean-field underdamped Langevin dynamics, Chen et al. (2023) establish its exponential rate of convergence and propose a finite-particle system that achieves a similar rate under assumptions that are easy to verify for the problem of training mean-field neural networks. In addition, Chen et al. (2023) implement an Euler-Maruyama discretization of the finite-particle system and show that it performs empirically faster when compared with the spacetime discretization of the mean-field Langevin dynamics in training a toy neural network model. However, spacetime discretization of the mean-field underdamped Langevin dynamics are not yet theoretically well understood. Furthermore, the rate obtained by Chen et al. (2023) for the dynamics does not resemble an accelerated rate when compared with Nitanda et al. (2022) and Chen et al. (2022) who analyze the overdamped setting.
@@
A remaining question is whether we can theoretically characterize the behavior of an implementable algorithm based on discretizing the mean-field underdamped dynamics. If there is a limiting bias, how does it scale with the number of particles and other problem parameters? Ideally, this characterization would give a sharper rate of convergence than Suzuki et al. (2023)’s spacetime discretization of the mean-field Langevin dynamics, suggesting there might be an advantage to adding momentum in the mean-field setting (at least in the worst case). In this paper, we introduce a fast implementable algorithm for solving EMO problems based on the mean-field underdamped Langevin dynamics. We prove that our proposed algorithm converges to a small limiting bias under a set of assumptions that subsumes many problems of interest. In particular, our contributions are summarized as follows.
@@
    We sharpen the convergence guarantees for both the mean-field underdamped Langevin dynamics and its space-discretization under the same set of assumptions utilized by Chen et al. (2023) (formalized in Theorems 3.1 and 3.2 and summarized in Table 1).
@@
    We introduce a novel spacetime discretization of the mean-field underdamped Langevin dynamics which we call the N-particle underdamped Langevin algorithm (Algorithm 1). We show the global convergence of our discretization in terms of the total variation (TV) distance. Importantly, our results improve on Suzuki et al. (2023)’s analysis of the spacetime discretization of the mean-field Langevin dynamics, and the additional assumptions needed to prove our result are satisfied by the same examples introduced by Suzuki et al. (2023); these results provide theoretical characterizations of momentum methods in several real-world applications including training neural networks, density estimation and sampling.
@@
The remainder of this work is organized as follows. Section 2 presents the formal definitions and assumptions as well as important related work. Section 3 proposes our main methods and theoretical results. Section 4 discusses the application of our methods to some classical problems. Section 5 describes our numerical experiments verifying the effectiveness of our proposed methods.
@@ 2.2 Background
We consider entropy-regularized mean-field optimization (EMO) problems, described by
where $\lambda>0$ and $F$ is a non-linear functional and $\text{Ent}(\mu)=\int\mu\log\mu$ is the negative entropy of $\mu$. The gradient flow structure in 2-Wasserstein metric of the EMO objective is called the mean-field Langevin dynamics,
@@
Hu et al. (2019) show that MLD converges to the solution of EMO problem, $\mu_{*}(x)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu_{*},x)\right),$ under mild conditions. We study the mean-field underdamped Langevin dynamics,
@@
and show a new sharp mixing-time bound; here, $\gamma>0$ is the damping coefficient; $\mu_{t}(x,v)=\text{Law}(x_{t},v_{t})$ and $\mu_{t}^{x}\triangleq\text{Law}(x_{t})=\int\mu_{t}(x,v)\mathrm{d}v$ is the $x$-marginal of $\mu_{t}$. The limiting distribution of MULD is the solution to the augmented EMO problem,
@@
where a momentum term is added to the EMO objective. To obtain the solution of the EMO problem, it suffices to marginalize the minimizer of (1) given by, $\mu_{*}(x,v)\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2\lambda}\|v\|^{2}\right).$ This work also sharpens the analysis of the space-discretization of MULD introduced by Chen et al. (2023), which we refer to as the $N$-particle underdamped Langevin dynamics,
@@
$\mathrm{d}v_{t}^{i}$ $=-\gamma v_{t}^{i}\mathrm{d}t\vspace{-10cm}-D_{\mu}F(\mu_{\textbf{x}_{t}},x^{i}_{t})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}^{i}_{t};$
$i=1,...,N$ where $\mu_{\textbf{x}}\triangleq\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$; $\mu^{i}_{t}\triangleq\text{Law}(x_{t}^{i},v_{t}^{i})$ and $(\mathrm{B}_{t}^{i})_{i=1}^{N}$ are $d$-dimensional Brownian motions independent of $i$. Note, a time-discretization of NULD is necessary to run the method on a machine. We explore two time-discretization techniques in this paper.
To motivate our algorithm as a time-discretization of NULD, we review discretizations of the underdamped Langevin dynamics (ULD), which is a special case of MULD where $F(\mu)=\int V(x)\mu(\mathrm{d}x)$ is a linear functional of $\mu$:
@@
for stepsize $h$, $\xi_{k}\sim\mathcal{N}(0,I_{d})$ and $t\in[kh,(k+1)h]$, has been well-studied and shown to incur the largest error (Kloeden et al., 1995; Platen and Bruti-Liberati, 2010). Recently, however, several works derive the convergence of a new ULMC algorithm in Wasserstein distance (Cheng et al., 2018), KL divergence (Ma et al., 2021) and Rényi divergence (Zhang et al., 2023) using a more precise discretization scheme, often referred to as the left point method (LPM) (Hu et al., 2021) and given by
@@
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-\nabla V(x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}_{t},$
@@ 2.3 Definitions and assumptions
Note that if $\hat{\mu}^{x}(x)\propto\exp(-\frac{\delta F}{\delta\mu}(\mu^{x},x))$ satisfies LSI with constant $\rho>0$, Assumption 2.3 is satisfied with the choice $\mathscr{C}_{\textsf{LSI}}=\frac{1}{2}\land\rho$. And we refer our readers to Chen et al. (2022, 2023); Suzuki et al. (2023) for the verification of Assumptions 2.1 and 2.3 in a variety of settings. Suzuki et al. (2023) consider a weaker smoothness assumption than Assumption 2.2 where they use $W_{2}$ distance. They verify $W_{2}$ smoothness for three examples including training mean-field neural networks, MMD minimization and KSD minimization, whereas Chen et al. (2022) verify $W_{1}$ smoothness only for training mean-field neural networks. In this paper, we verify $W_{1}$ smoothness in Assumption 2.2 for the other two examples. Beyond Assumptions 2.1-2.3, we introduce four additional assumptions that are sufficient for our spacetime discretization analysis.
@@
Notably, Suzuki et al. (2023) assume that $F$ can be decomposed as $F(\mu)=U(\mu)+\mathbb{E}_{x\sim\mu}[r(x)]$ where $\|D_{\mu}U(\mu,x)\|\leq R$ for any $\mu\in\mathcal{P}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, and where $r(x)$ is $\lambda_{2}$-smooth with $\nabla r(0)=0$ in order to establish the convergence of their spacetime discretization of MLD. Thus, their assumption, $\|D_{\mu}F(\mu,x)\|\leq\|D_{\mu}U(\mu,x)\|+\|\nabla r(x)\|\leq R+\lambda_{2}\|x\|$, implies Assumption 2.4 with the choice $\mathscr{L}\geq\max\{R,\lambda_{2}\}$. The next three assumptions are needed for bounding the second moment of the iterates $(x_{t},v_{t})$ and $(x_{t}^{i},v_{t}^{i})$ along MULD and NULD, which is crucial for the establishment of our discrete-time convergence:
@@ 3 N-particle underdamped Langevin algorithm
$\mathrm{d}v_{t}$ $=-\gamma v_{t}\mathrm{d}t-D_{\mu}F(\mu^{x}_{k},x_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}_{t},$
@@
$\mathrm{d}v_{t}^{i}$ $=\gamma v^{i}_{t}\mathrm{d}t-D_{\mu}F(\mu_{\textbf{x}_{k}},x^{i}_{k})\mathrm{d}t+\sqrt{2\gamma}\mathrm{dB}^{i}_{t}$
@@ 3.1 Convergence analysis
spacetime discretizations NLA (Suzuki et al., 2023) ${\Theta}\left(\kappa\mathscr{L}^{3}d/\epsilon^{2}\right)$ $\widetilde{\Theta}\left(\kappa^{2}\mathscr{L}d/\epsilon^{2}\right)$
@@ 4 Applications of Algorithm 1
In Section 1, we briefly mention three examples on the applications of problem (2.2). In this section, we will show how Algorithm 1 can be applied to these examples by verifying Assumptions 2.1-2.7 for these examples. We defer the proof of verification to Appendix B.